OpenStudy (anonymous):

What is the 8th term of the geometric sequence where a1 = 256 and a3 = 16?

OpenStudy (lilg132):

a = 256 a + 2d = 16 d = -120 a + 7d = 257 + 7(-120) = 257 + 840 = 1097 is 1097 one of your answers?

OpenStudy (anonymous):

nahh, i got 0.015625

OpenStudy (lilg132):

oops i got my negative mixed up

OpenStudy (lilg132):

-583?

OpenStudy (anonymous):

nahh, it was 0.015625, thank you for your help though (:

OpenStudy (lilg132):

hmm thats strange its ok